A 60.0 kg sprinter starting a 100.0 m dash accelerates at a rate of 3.0 m/s2. What is the force of the sprinter?


1
Expert's answer
2021-02-15T00:53:48-0500

Given quantities:

  m = 60.0 [kg]              S = 100.0 [m]                a = 3.0 ["\large\frac{m}{s^2}"]

The force of sprinter is found by Newton's second law - furthermore

"F = ma = 60.0*3 \space kg\large\frac{m}{s^2}" "= 180N"